# Towards Operational Automated Greenhouse Gas Plume Detection

### v1

@@ 4.4 Models
GoogLeNet: The GoogLeNet architecture (Szegedy et al., 2015) incorporates several “inception blocks” designed to learn convolutional filters capturing discriminative image patterns at multiple spatial scales. Designed for image classification tasks, we train GoogLeNet using tiles extracted from CMF image scenes with tilewise plume or background labels determined by the presence of plume pixels in each tile. We compare the capabilities of GoogLeNet-based plume image classifier models to those of the recently proposed MethaNet (Jongaramrungruang et al., 2022) plume image classifier model on a variety of plume image classification experiments (see Sections LABEL:apx:bias_region and LABEL:apx:bias_distrib). To perform pixelwise plume segmentation, we convert our trained GoogLeNet image classifier to a “shift and stitch” Fully Convolutional Network (FCN) via the procedure described in Shelhamer et al. (2015). This allows us to efficiently apply GoogLeNet image classifiers to full-scene, pixelwise plume segmentation and instance detection tasks. We also apply model anti-aliasing as described in Zhang (2019) to avoid shift invariance artifacts in produced masks (Lee et al., 2020).
@@
U-Net: The U-net architecture Ronneberger et al. (2015) is among the most widely-used CNN architectures to date. Originally designed for medical applications involving pixelwise ”semantic segmentation” tasks Siddique et al. (2021), U-net is the base plume detector model used in several recent papers (Bruno et al., 2023; Růžička et al., 2023). The U-net architecture learns detailed image patterns directly from pixelwise labeled training images by connecting each convolution/downsampling layer to a matching deconvolution/upsampling layer. These ‘skip connections” allow the model to learn discriminative image filters at multiple spatial scales (via convolution & pooling) while simultaneously learning to generate accurate pixelwise predictions at native spatial resolution by combining the outputs of learned multiscale filters.
@@
Tilewise Model: “Shift & Stitch” GoogLeNet FCN. Since this model was trained for image classification tasks using tilewise (rather than pixelwise) labels, it offers advantages for plume instance detection but is ill-suited for pixelwise segmentation due to very coarse masks.
@@
Pixelwise Model: Standard U-net optimized with purely pixelwise objective. Generates detailed pixelwise predictions but is ill-suited for instance detection due to a large number of false positives.
@@
Multitask Model: Identical U-net model as pixelwise model optimized with multitask segmentation & classification objective. Provides a compromise between plume instance detection and pixelwise segmentation objectives.
@@ 5 Results
5.1 Scenewise Plume Instance Detection and Segmentation
@@
Comparing AVIRIS-NG Plume Detector Models: While the tilewise-trained model yields the highest instance detection accuracy overall, it performs poorly pixelwise due to the coarse segments it generates. Consequently, excluding small/weak detection candidates produces the smallest increases in both pixelwise and instance-level accuracy of the three models.
The pixelwise-trained model gives a slight improvement in pixelwise accuracy, but produces the worst overall instance detection accuracy due to the quantity of small false positive detections it generates. Consequently, excluding ambiguous (small/weak) detection candidates produced by the pixelwise model provides the most substantial increase in detection accuracy of the three models.
The multitask model is less prone to overfitting to ambiguous/uncertain plume boundary pixels in the training data, which accounts for its increase in pixelwise accuracy over the pixelwise-trained model. It also performs competitively with the tilewise model in terms of instance detection, particularly when we exclude small/weak ambiguous detections.
Comparing AVIRIS-NG and EMIT Results: With respect to pixelwise segmentation performance, AVIRIS-NG pixelwise mispredictions are dominated by false positives (precision recall) while false negatives dominate the EMIT pixelwise mispredictions (precision recall). Several factors contribute to the mismatched pixelwise performance of the AVIRIS-NG and EMIT models on their respective validation scenes.
The primary difference lies in the methodology used to define pixelwise plume labels for the airborne platforms versus the EMIT platform; EMIT plume labels are inclusive by design, while AVIRIS-NG plume labels are intentionally conservative. EMIT plume labelers are instructed to include a small pixel buffer around any plume they annotate, so the labels for each EMIT plume instance will capture some background pixels outside of each plume’s boundary. These low-concentration pixels account for the higher pixelwise false negative rate (lower recall) on the EMIT plumes. In contrast, the cmf-guided AVIRIS-NG plume labels only include high-concentration pixels near each plume’s origin point that exceed our 500 background enhancement threshold. As a result, the pixelwise labels of AVIRIS-NG plume instances often do not capture the full spatial extent of each plume, often excluding boundary pixels and “disconnected” plume components. These unlabeled pixels account for the higher pixelwise false positive rate (lower precision) on the AVIRIS-NG validation scenes. Additionally, the EMIT VISIONS data set only includes high SnR CMFs approved by multiple reviewers, while the AVIRIS-NG validation scenes were inspected and reviewed by fewer reviewers overall.
While differences in plume labeling methodologies broadly account for the mismatched distributions of pixelwise prediction errors between AVIRIS-NG and EMIT results, sensor and retrieval-specific differences help explain the gap in pixelwise precision between the AVIRIS-NG and EMIT results. EMIT is a more efficient sensor than AVIRIS-NG, and CMFs derived from EMIT data use an improved, full-VSWIR-based retrieval compared to the original SWIR2-only retrieval approach applied to AVIRIS-NG data from early airborne imaging campaigns. Compared to the AVIRIS-NG CMF products, plume and background pixels tend to be better separated in terms of local pixelwise contrast, and fewer false enhancements are present in EMIT CMFs. This is also reflected in the EMIT instance detection performance, where the multitask model performs better in both precision and recall.
@@ 5.2 Evaluating Regional Generalization
We evaluate generalization performance across distinct regions observed in AVIRIS-NG and GAO airborne imaging campaigns. However, since we do not have scenewise quality-controlled data for the GAO campaigns, we restrict our evaluation to tilewise instance classification performance results based on the Tilewise-trained model in this section.
Figure 8 summarizes the tilewise prediction accuracy of the AVIRIS-NG multicampaign model applied to the AVIRIS-NG training/test tiles and the GAO test tiles detailed in Table 5. In all cases the FPR dominates—typically by a factor of 2-3 the FNR, indicating that false positives constitute most of the prediction errors.
Figure 9 shows example false positive mispredictions produced by the AVIRIS-NG multicampaign model. All are false enhancements with concentration levels consistent with real plumes, though their morphologies vary according to the the phenomena that produce them. For instance, linear false enhancements that follow the downtrack imaging direction are MF artifacts produced when outliers corrupt the background covariance matrix used by the retrieval algorithm, while other linear false enhancements are caused by confuser materials (e.g., agricultural plastics, fertilizer, rooftop, or road surfaces with absorptions consistent with some hydrocarbons) coupled with sharp local changes in albedo.
@@
6 Measuring Airborne Plume Detectability
Having shown that the multicampaign model yields satisfactory results when applied to observations from similar regions, we now turn to the task of measuring plume detectability based on the characteristics of the labeled test plumes identified in those regions. As discussed in Section 2.5, we do not consider instances where enhancements fall below the range of instrument detectability. In the context of plumes observed by airborne spectrometers, our analysis below aims to address the following questions:
1.
How do plume size and concentration impact plume detectability?
2.
Can we reliably detect plumes generated by some classes of emission sources more reliably than others?
We base our assessment on the plumes observed in all airborne campaigns excluding those from the AVIRIS-NG Four Corners and GAO Penn campaigns. We exclude these due to their unique challenges and characteristics described in Suppl. Section LABEL:apx:bias_region.
Figure 11 gives the tilewise FNR scores for all AVIRIS-NG and GAO test plumes listed in Table 5 with respect to the average concentration of their constituent pixels (x-axis) versus the total area of each plume (y-axis). Most mispredictions occur at low average pixelwise concentrations () where plume and background enhancements are numerically similar. Weak plumes with small spatial footprints (area ) are the most commonly mispredicted (FNR=34.6%) since they are difficult to distinguish from false enhancements of similar size—such as the high albedo false enhancements shown in 9.
Figure 13 gives the average FNR (top) and the average number of false negatives (bottom) per-campaign with respect to the source emission sectors of the identified plumes in the selected airborne AVIRIS-NG and GAO campaigns.
Figure 12 gives the percentage of sources observed in each campaign and the distribution of source sectors for sources observed across those campaigns.
We observe a significantly higher false negative rate for plumes associated with Electricity Generation (EG) sources – roughly higher on training data, and nearly higher on test data – relative to the other sectors. The issues with detecting plumes from EG sources are the result of two issues. First, all of the plumes we observed from EG sources occur in urban areas with diverse surface materials and substantial albedo variation. As shown in Figure 1, CMF images representing complex and diverse urban areas often contain similarly complex false enhancements, with many occurring in the relative vicinity of observed plumes. Second, EG sources only represent 1.3% of the all of the sources observed in the selected airborne campaigns. Given the complexity of the CMF enhancements where EG sources occur, it is unlikely we have a representative sample of plumes produced by EG sources to train our model to detect them robustly. Interestingly, we observe the lowest FNR of all sectors on wastewater treatment facilities (% of all observed airborne sources) which also occur in similarly diverse urban areas, but are also likely driven by the lack of representative samples. In terms of the overall quantity of false negatives per campaign, oil and natural gas (Oil&NG) sources dominate by a factor of on training plumes, and nearly on test plumes, relative to the average false negative count of the remaining source sectors. This is not particularly surprising given that Oil&NG sources represent 86% of the all of the identified sources observed in the airborne campaigns we consider in this work. A less obvious contributing factor to the high Oil&NG FNR count is that many false negatives represent Oil&NG plumes observed in the AVIRIS-NG and GAO Permian campaigns with small pixelwise footprints. Additionally, observations from the AVIRIS-NG Permian campaign are lower resolution than those from other AVIRIS-NG campaigns (7.4 vs. 3.1 GSD), and observations from the GAO Permian campaign are lower resolution (6.0 vs. 4.67 GSD) than the other GAO campaigns.
The relative impact of plume size and concentration on plume detectability vary according to the capabilities of plume detector model, spatial resolution, sensor platform, and retrieval approach. While challenges involving the detection of specific emission source types are broadly relevant, detectability tradeoffs between plume area and concentration level are sensor and data specific, and our results shown in 11 are only directly applicable to AVIRIS-NG & GAO CMF products derived using the SWIR2-only retrieval from fine-resolution observations with GSD 3-7. That said, we can reasonably expect that replicating our AVIRIS-NG and GAO-based experiments with analogous data captured by advanced airborne sensors like AVIRIS-3, or replacing the SWIR2 CMF retrievals with full-VSWIR CMF retrievals in the experiments above will improve detectability at all plume sizes and concentrations. However, the relative impact of plume concentration and area on plume detectability are less clear for the full-VSWIR retrievals derived from coarse resolution spaceborne observations captured by spaceborne platforms like EMIT and the Carbon Mapper Coalition’s Tanager-1. Despite their advanced sensing capabilities relative to analogous airborne sensors like AVIRIS-NG and GAO, subpixel averaging in coarse resolution spaceborne radiance data at average GSD of 60 inevitably reduces retrieved concentration levels relative to retrievals derived using fine resolution airborne spectrometer data. Consequently, small or low concentration plumes that are detectable in airborne spectrometer observations are often undetectable (or unidentifiable) in spaceborne observations.

### v2

@@ 4.4 Models
GoogLeNet: The GoogLeNet architecture (Szegedy et al., 2015) incorporates several “inception blocks” designed to learn convolutional filters capturing discriminative image patterns at multiple spatial scales. Designed for image classification tasks, we train GoogLeNet using tiles extracted from CMF image scenes with tilewise plume or background labels determined by the presence of plume pixels in each tile. We compare the capabilities of GoogLeNet-based plume image classifier models to those of the recently proposed MethaNet (Jongaramrungruang et al., 2022) plume image classifier model on a variety of plume image classification experiments (see Sections LABEL:apx:bias_region and LABEL:apx:bias_distrib). To perform pixelwise plume segmentation, we follow the procedure described by Long et al. (2015) to convert our trained GoogLeNet image classifier to a Fully Convolutional Network (FCN) by replacing the fully connected layers with convolutional filters that retain the learned parameters. To obtain a dense, high-resolution output, we then employ the “shift and stitch” procedure to account for the spatial downsampling inherent in the pooling layers. This allows us to efficiently apply GoogLeNet image classifiers to full-scene, pixelwise plume segmentation and instance detection tasks. We also apply model anti-aliasing as described in Zhang (2019) to avoid shift invariance artifacts in produced masks (Lee et al., 2020).
@@
U-Net: The U-net architecture Ronneberger et al. (2015) is among the most widely-used CNN architectures to date. Originally designed for medical applications involving pixelwise ”semantic segmentation” tasks Siddique et al. (2021), U-net is the base plume detector model used in several recent papers (Bruno et al., 2024; Růžička et al., 2023). The U-net architecture learns detailed image patterns directly from pixelwise labeled training images by connecting each convolution/downsampling layer to a matching deconvolution/upsampling layer. These ‘skip connections” allow the model to learn discriminative image filters at multiple spatial scales (via convolution & pooling) while simultaneously learning to generate accurate pixelwise predictions at native spatial resolution by combining the outputs of learned multiscale filters.
@@
Tilewise Model: The GoogLeNet FCN model architecture with “shift-and-stitch” initially trained for the plume classification task using tilewise (rather than pixelwise) labels. Due to its optimization objective, it offers advantages for plume instance detection but is ill-suited for pixelwise segmentation as it produces relatively coarse and imprecise masks. Complete details about this model and its optimization objective is provided in Suppl. Section LABEL:sub:tilewise.
@@
Pixelwise Model: The U-Net model architecture trained for the plume segmentation task optimized with purely pixelwise optimization objective. It generates detailed pixelwise saliency map predictions, but it is ill-suited for instance detection due to a large number of false positives. Complete details about this model and its optimization objective is provided in Suppl. Section LABEL:sub:pixelwise.
@@
Multitask Model: The U-net model architecture trained for the plume segmentation task optimized with both segmentation and classification task objectives (hence, multitask). Provides a compromise with good instance detection performance (reduced false positives) and detailed saliency map predictions. Complete details about this model and its optimization objective is provided in Suppl. Section LABEL:sub:multitask.
@@ 5 Results
5.1 Quantitative Scenewise Detection and Segmentation Performance
In this section, we describe the scenewise detection and segmentation performances of our airborne and spaceborne models in detail. As described in Section 3 and further in Suppl. Section LABEL:sub:instance, we calculate our metrics on entire flightlines and scenes, not only the chipped tiles that the models were trained on. By evaluating both detection and segmentation metrics, we aim to characterize the models’ performances in contexts relevant to operational use.
5.1.1 Tradeoffs Between AVIRIS-NG Models
Figure 5 compares the pixelwise segmentation and instance detection performances of the three AVIRIS-NG models, and Figure 6 visualizes examples of detected plumes and their segmentations. The tilewise model, only optimized for classification, yields the highest instance detection performance (0.64 precision and 0.71 recall). However, partially because it was not trained on any label masks, it generates relatively coarse masks, as shown in the second column of Figure 6 (0.13 precision and 0.68 recall). Consequently, excluding ambiguous small/weak detection candidates (indicated in light blue and yellow) produces the smallest increases in both instance-level (+0.05 precision and +0.04 recall) and pixelwise accuracy (+0.02 precision and +0 recall) of the three models.
The pixelwise-trained model gives a slight improvement in pixelwise accuracy (0.17 precision and 0.79 recall), but produces the worst overall instance detection accuracy due to the large quantity of small false positive detections it generates (0.18 precision and 0.77 recall). Consequently, excluding ambiguous small/weak candidates produced by the pixelwise model provides the most substantial increase in detection accuracy of the three models by filtering out spurious false positive detections (+0.11 precision and +0.03 recall). However, it is not enough to sufficiently improve the instance detection performance for operational use.
The multitask model, optimized for both classification and segmentation, is less prone to overfitting to ambiguous/uncertain plume boundary pixels in the training data, which accounts for its increase in pixelwise accuracy (0.35 precision, 0.63 recall) over the pixelwise-trained model. It also performs competitively with the tilewise model in terms of instance detection (0.49 precision, 0.66 recall), particularly when we exclude small/weak ambiguous detections (+0.11 precision, +0 recall). By accounting for both detection and segmentation in its optimization, the multitask model provides the best tradeoff between detecting the presence of plumes and producing detailed masks.
5.1.2 Comparing AVIRIS-NG and EMIT Results
Figure 5 also compares the pixelwise segmentation and instance detection performances of the multitask EMIT model. With respect to pixelwise segmentation performance, AVIRIS-NG pixelwise mispredictions are dominated by false positives (0.35 precision 0.63 recall) while false negatives dominate the EMIT pixelwise mispredictions (0.71 precision 0.52 recall). Several factors contribute to the mismatched pixelwise performance of the AVIRIS-NG and EMIT models on their respective validation scenes.
The primary difference lies in the methodology used to define pixelwise plume labels for the airborne platforms versus the EMIT platform. As described in Section 2.3, EMIT plume labels are inclusive by design, while AVIRIS-NG plume labels are intentionally conservative. EMIT plume labelers are instructed to include a small pixel buffer around any plume they annotate, so the labels for each EMIT plume instance will capture some background pixels outside of each plume’s boundary. These low-concentration pixels account for the lower recall (0.52; higher false negative rate) on the EMIT plumes. In contrast, the CMF-guided AVIRIS-NG plume labels only include high-concentration pixels near each plume’s origin point that exceed our 500ppm-m background enhancement threshold. As a result, the pixelwise labels of AVIRIS-NG plume instances often do not capture the full spatial extent of each plume, often excluding boundary pixels and “disconnected” plume components. These unlabeled pixels account for the lower pixelwise precision (0.35; higher false positive rate) on the AVIRIS-NG validation scenes. Additionally, the EMIT VISIONS data set only includes high SnR CMFs approved by multiple reviewers, while the AVIRIS-NG validation scenes were inspected and reviewed by fewer reviewers overall.
While differences in plume labeling methodologies broadly account for the mismatched distributions of pixelwise error between AVIRIS-NG and EMIT results, sensor and retrieval-specific differences also help explain the gap in pixelwise precision between the AVIRIS-NG and EMIT results (0.35 and 0.71, respectively). EMIT is a more efficient sensor than AVIRIS-NG, and CMFs derived from EMIT data use an improved, full-VSWIR-based retrieval compared to the original SWIR2-only retrieval approach applied to AVIRIS-NG data from early airborne imaging campaigns. Compared to the AVIRIS-NG CMF products, plume and background pixels tend to be better separated in terms of local pixelwise contrast, and fewer false enhancements are present in EMIT CMFs. This is also reflected in the EMIT instance detection performance, where the multitask model performs better in both recall (0.59) and precision (0.82).
5.2 Salience Map Comparison Between Models
@@ 5.3 Evaluating Regional Generalization
We evaluate the generalization performance of the AVIRIS-NG multicampaign model across distinct regions observed in AVIRIS-NG and GAO airborne imaging campaigns. However, since we do not have scenewise quality-controlled data for the GAO campaigns, we restrict our evaluation to tilewise instance classification performance results. Figure 7 summarizes the tilewise classification accuracy of the AVIRIS-NG multicampaign model applied to the AVIRIS-NG train/test tiles and the GAO test tiles detailed in Table 1.
In all cases, the lower precision indicates that the false positives constitute most of the prediction errors. Still, the multicampaign model yields similar results on the AVIRIS-NG Cal, Permian, and COVID train/test datasets, as expected. On the AVIRIS-NG Four Corners test dataset, the multicampaign model performs similarly to the other AVIRIS-NG datasets, as the Four Corners region is geographically similar to California and the Texas Permian Basin. This is also reflected in the results for the GAO Permian and CARB (California), and Denver (Colorado) datasets.
However, the model has a significantly reduced precision on the GAO Penn (Pennsylvania) dataset, which is geographically distinct from the rest of the regions. This confirms that the regional generalization of the AVIRIS-NG multicampaign model is limited to regions that share the geographical characteristics of California and Texas. While not demonstrated here, we would expect the EMIT model trained on globally sampled plumes and background scenes to have further improved global generalizability. We also present experiments where we train and evaluate models trained on individual campaigns in Suppl. Section LABEL:apx:bias_region, demonstrating that training on multiple campaigns is necessary for regional generalization.
Figure 8 shows example false positive mispredictions produced by the AVIRIS-NG multicampaign model. All are false enhancements with concentration levels consistent with real plumes, though their morphologies vary according to the the phenomena that produce them. For instance, linear false enhancements that follow the downtrack imaging direction are CMF artifacts produced when outliers corrupt the background covariance matrix used by the retrieval algorithm, while other linear false enhancements are caused by confuser materials (e.g., agricultural plastics, fertilizer, rooftop, or road surfaces with absorptions consistent with some hydrocarbons) coupled with sharp local changes in albedo.
@@
5.4 Measuring Airborne Plume Detectability
We evaluate the limits of our models’ plume detectability based on the physical characteristics and source sectors of the labeled plumes in our test dataset. As discussed in Section 2.5, we do not consider instances where enhancements fall below the range of instrument detectability. Of course, detectability limits are intrinsically tied to the spatial resolution and sensitivity of the instrument, platform, and retrieval, not only the model performance. In this section, we focus on plumes observed by the AVIRIS-NG and GAO airborne campaigns.
5.4.1 Impact of Physical Characteristics on Detectability
We begin by assessing the positive and negative detections of plumes from the AVIRIS-NG and GAO test datasets by the AVIRIS-NG multicampaign multitask model. We exclude the AVIRIS-NG Four Corners and GAO Penn datasets due to anomalies in data quality further described in Suppl. Section LABEL:apx:bias_region. We characterize the physical characteristics of each plume by determining its area of delineation and the average concentration within that area. Sorting the plumes by these two properties allows us to highlight correlations with model performance. For the purposes of this analysis, we use the tilewise False Negative Rate (FNR), as each tile is cropped around each plume.
Figure 10 shows these tilewise FNR scores calculated for binned ranges of plume area (y-axis) and average concentration (x-axis). As expected, most mispredictions occur at low average pixelwise concentrations ( ) where plume enhancements are less distinguishable from background enhancements. Further, such weak concentration plumes with also small areas ( ) are the most commonly mispredicted, with a false negative rate of 34.6%. These are especially difficult to distinguish from false enhancements of similar size, as there are less distinct differences in spatial morphology—such as the high albedo false enhancements shown in Figure 8. By contrast, plumes with similarly small area but higher concentration (25 to 50 ) are detected far more reliably (8.6% FNR), whereas plumes with similarly small concentration but larger area (2.5 to 10.0 ) are still missed often (31.2% FNR). Despite the larger plume area, a plume with concentrations that cannot be reliabily distinguished from the background cannot be detected.
5.4.2 Impact of Source Sector Imbalance
In addition to identifying plumes in the AVIRIS-NG and GAO campaign datasets, human experts also determined the IPCC sectors of each plume based on identified ground infrastructure. The distribution of source sectors (coal mining, electricity generation, livestock, oil & gas, solid waste, and waste water) for sources observed across these campaigns is shown in Figure 11. As expected from the US regions observed, oil & gas sources dominate the dataset (86%), with few sources attributed to solid waste (6%) and livestock (5%), and minimal representation of coal mining (1.6%), electricity generation (1.3%), and waste water (0.2%). In such cases of high sector imbalance, we expect data-driven machine learning models to fail on severely underrepresented classes, as there are not enough examples to learn from. We also base this on the fact that plumes from different sectors have different observed characteristics; for example, oil & gas plumes from flare stacks are at higher pressure and temperature than livestock plumes from manure management systems.
We investigate these effects by calculating the average False Negative Rate across campaigns for each source sector. We again exclude the AVIRIS-NG Four Corners and GAO Penn datasets, which also excludes the coal mining sector (mostly represented by the GAO Penn campaign). In this analysis, we calculate the FNR for both the training and test sets. A perfect machine learning model would have a 0% FNR on any plumes in the training dataset (as it will have learned from them), but in practice single-digit FNRs are expected.
Figure 12 shows the average FNR (top) and the average number of false negatives (bottom) across campaigns with respect to the source emission sectors of the identified plumes. For oil & gas, a well-represented sector (86%), the training set FNR is very low and the test set FNR is slightly higher, as expected. The average number of false negative detections dominates that of other sectors, only because it already represents the vast majority of all identified sources. Further analysis of these false negative detections reveals that many of these plumes have very small areas, which are difficult to detect as discussed previously. For lesser-represented sectors, livestock (5%) and solid waste (6%), both training set and test set FNRs remain below 10%, indicating there was sufficient representation in the training set for the model to learn how to detect these plumes. It is also possible that these plumes shared similar morphological features as plumes attributed to other sectors.
We observe a significantly higher false negative rate for plumes associated with Electricity Generation (EG) sources—0.22 on the training set and 0.33 on the test set—relative to the other sectors. This indicates that 1.3% of the dataset distribution was not sufficient for the model to learn to detect plumes attributed to EG. This effect is further exacerbated by the fact that all of the plumes we observed from EG sources occur in urban areas with diverse surface materials and substantial albedo variation. As shown in Figure 1, CMF images representing complex and diverse urban areas often contain similarly complex false enhancements, with many occurring in the relative vicinity of observed plumes. A more representative sample of EG plumes were necessary for the model to learn to detect them robustly.
Interestingly, we observe the lowest FNR (0%) of all sectors on wastewater treatment facilities (0.2% of all observed airborne sources) which also occur in similarly diverse urban areas, but are also likely driven by the lack of representative samples. Rather than assigning a systematic justification, we conclude these plumes may have been sufficiently high concentration to be detected regardless of sector-correlated factors, or the model may have been lucky on the few plumes included in the evaluation.